OpenStudy (anonymous):

Solve for x. x ^2 + 8x + 16 = 25

OpenStudy (anonymous):

1

OpenStudy (anonymous):

there's 2 answers

OpenStudy (anonymous):

1,-9

OpenStudy (anonymous):

x^2 - 8x + 16 = 25 (factor left side as perfect square trinomial) (x - 4)^2 = 25 (square root both sides) x - 4 = ±5 (add 4 to both sides) x = 4 ± 5 x = 4 + 5 or y = 4 - 5 x = 9 or y = -1 <===You answer :) Hope it helps!

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

your welcome!